# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ 2.2 Linear Function Approximation and Linear TD Learning
One can check that the linear projected Bellman operator ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ is a $\gamma$-contraction in the Polish space $({\mathscr{V}}_{{\bm{\phi}}},\left\|{\cdot}\right\|_{\mu_{\pi}})$. Hence, ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ admits a unique fixed point ${\bm{V}}_{{\bm{\psi}}^{\star}}$, which satisfies $\|{\bm{V}}^{\pi}-{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}\leq(1-\gamma^{2})^{-1/2}\|{\bm{V}}^{\pi}-{\bm{\Pi}}_{{\bm{\phi}}}^{\pi}{\bm{V}}^{\pi}\|_{\mu_{\pi}}$ (Theorem 9.8 Bellemare et al., 2023). In Appendix B.1, we show that ${\bm{\psi}}^{\star}\in{\mathbb{R}}^{d}$ is the unique solution to the linear system for ${\bm{\psi}}\in{\mathbb{R}}^{d}$:
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right].$ (2)
@@
In the subscript of the expectation, we abbreviate $s\sim\mu_{\pi}(\cdot),a\sim\pi(\cdot|s),r\sim{\mathcal{P}}_{R}(\cdot|s,a),s^{\prime}\sim P(\cdot|s,a)$ as $s,a,r,s^{\prime}$. For brevity, we will continue to use such abbreviations in this paper when there is no ambiguity. We can use LSA to solve the linear projected Bellman equation (Eqn. (2)). As a result, at the $t$-th time-step, the updating scheme of Linear-TD is
${\bm{\psi}}_{t}={\bm{\psi}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left[{\left({{\bm{\phi}}(s_{t})-\gamma{\bm{\phi}}(s_{t+1})}\right)^{\intercal}{\bm{\psi}}_{t-1}-r_{t}}\right].$ (3)
See Appendix B.2 for convergence results for Linear-TD.
2.3 Distributional Policy Evaluation and Distributional TD Learning
In certain applications, we are not only interested in finding the expectation of the random return $G^{\pi}(s)$ but want to find the whole distribution of $G^{\pi}(s)$. This task is called distributional policy evaluation. We use $\eta^{\pi}(s)\in{\mathscr{P}}$ to denote the distribution of $G^{\pi}(s)$ and let ${\bm{\eta}}^{\pi}:=(\eta^{\pi}(s))_{s\in{\mathcal{S}}}\in{\mathscr{P}}^{\mathcal{S}}$. Then ${\bm{\eta}}^{\pi}$ satisfies the distributional Bellman equation:
$\eta^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),r\sim{\mathcal{P}}_{R}({\cdot}\mid s,a),s^{\prime}\sim P(\cdot\mid s,a)}[\left({b_{r,\gamma}}\right)_{\#}\eta^{\pi}(s^{\prime})],\quad\forall s\in{\mathcal{S}}.$ (4)
Here $b_{r,\gamma}\colon{\mathbb{R}}\to{\mathbb{R}}$ is the affine function defined by $b_{r,\gamma}(x)=r+\gamma x$, and $f_{\#}\nu$ is the push forward measure of $\nu$ through any function $f\colon{\mathbb{R}}\to{\mathbb{R}}$, so that $f_{\#}\nu(A)=\nu(f^{-1}(A))$ for any Borel set $A$, where $f^{-1}(A):=\left\{{x\colon f(x)\in A}\right\}$. The distributional Bellman equation can also be written as ${\bm{\eta}}^{\pi}={\bm{\bm{\mathcal{T}}}}^{\pi}{\bm{\eta}}^{\pi}$. The operator ${{\bm{\mathcal{T}}}}^{\pi}\colon{\mathscr{P}}^{\mathcal{S}}\to{\mathscr{P}}^{\mathcal{S}}$ is called the distributional Bellman operator and ${\bm{\eta}}^{\pi}$ is the associated fixed point.
In analogy to TD learning, we can solve the distributional Bellman equation by LSA and get the distributional TD learning algorithm given the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$:
@@ 2.3 Distributional Policy Evaluation and Distributional TD Learning
$\eta_{t}(s_{t})=\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})].$
@@
We comment the algorithm above is not computationally feasible as we need to manipulate an infinite-dimensional object at each iteration.
2.4 Categorical Parametrization of Return Distributions and Categorical TD Learning
@@
In order to deal with the infinite-dimensional return distributions in a computational tractable manner, we consider the finite-dimensional categorical parametrization as in (Bellemare et al., 2017; Rowland et al., 2018; Bellemare et al., 2019; Rowland et al., 2024b; Peng et al., 2024). To incorporate function approximation in the distribution parametrization which cannot guarantee non-negative outputs, we will work with the signed measure space ${\mathscr{P}}^{\mathsf{sign}}$ with total mass $1$ instead of the standard probability space ${\mathscr{P}}\subset{\mathscr{P}}^{\mathsf{sign}}$:
@@
${\mathscr{P}}^{\mathsf{sign}}:=\left\{\nu\colon\left|{\nu}\right|({\mathbb{R}})<\infty,\nu({\mathbb{R}})=1,\operatorname{supp}(\nu)\subseteq\left[0,(1-\gamma)^{-1}\right]\right\}.$
@@
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, we define its cumulative distribution function (CDF) as $F_{\nu}(x):=\nu[0,x]$. We can naturally define the $L^{2}$ and $L^{1}$ distance between CDFs as Cramér distance $\ell_{2}$ and $1$-Wasserstein distance $W_{1}$ in ${\mathscr{P}}^{\mathsf{sign}}$ respectively. The distributional Bellman operator (see Eqn. (4)) can also be extended to the product space $({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$ without modifying its definition.
@@
${\mathscr{P}}^{\mathsf{sign}}_{K}:=\left\{\nu_{\bm{p}}=\sum_{k=0}^{K}p_{k}\delta_{x_{k}}\colon{\bm{p}}=\left({p_{0},\cdots,p_{K-1}}\right)^{\intercal}\in{\mathbb{R}}^{K},p_{K}=1-\sum_{k=0}^{K-1}p_{k}\right\},$
which is an affine subspace of ${\mathscr{P}}^{\mathsf{sign}}$, where $K\in{\mathbb{N}}$, $\left\{{x_{k}=k\iota_{K}}\right\}_{k=0}^{K}$ are equally-spaced points of the support, and $\iota_{K}=\left[{K(1-\gamma)}\right]^{-1}$ is the gap between two adjacent points. We say ${\bm{p}}$ is the probability mass function (PMF) representation of $\nu_{\bm{p}}\in{\mathscr{P}}^{\mathsf{sign}}_{K}$. In fact, this representation is an isometry (see Proposition B.1). We define the categorical projection operator ${\bm{\Pi}}_{K}\colon{\mathscr{P}}^{\mathsf{sign}}\to{\mathscr{P}}^{\mathsf{sign}}_{K}$ as
@@
Following (Proposition 5.14 Bellemare et al., 2023), one can show that $\bm{\Pi}_{K}\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ is uniquely identified with a vector ${\bm{p}}_{\nu}=\left({p_{k}(\nu)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, where
$p_{k}(\nu)={\mathbb{E}}_{X\sim\nu}[(1-\left|{(X-x_{k})/{\iota_{K}}}\right|)_{+}]:=\int_{\left[{0,(1-\gamma)^{-1}}\right]}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx).$ (5)
$\bm{\Pi}_{K}$ is in fact an orthogonal projection (see Proposition B.2), and thus is non-expansive.
For any ${\bm{\eta}}\in\left({{\mathscr{P}}^{\mathsf{sign}}}\right)^{\mathcal{S}}$, $s\in{\mathcal{S}}$, we define ${\bm{p}}_{\bm{\eta}}(s):={\bm{p}}_{\eta(s)}$ and lift ${\bm{\Pi}}_{K}$ to the product space by defining $\left[{\bm{\Pi}_{K}{\bm{\eta}}}\right](s):=\bm{\Pi}_{K}\eta(s)$. One can check that the categorical Bellman operator ${\bm{\Pi}}_{K}{\bm{\mathcal{T}}}^{\pi}$ (see Proposition B.3 for the characterization of $\bm{\Pi}_{K}\bm{\mathcal{T}}^{\pi}$) is a $\sqrt{\gamma}$-contraction in the Polish space $(({\mathscr{P}}_{K}^{\mathsf{sign}})^{\mathcal{S}},\ell_{2,\mu_{\pi}})$, where $\ell_{2,\mu_{\pi}}(\bm{\eta}_{1},\bm{\eta}_{2}):=({\mathbb{E}}_{s\sim\mu_{\pi}}[\ell_{2}^{2}(\eta_{1}(s),\eta_{2}(s))])^{1/2}$ is the $\mu_{\pi}$-weighted Cramér distance between $\bm{\eta}_{1},\bm{\eta}_{2}\in({\mathscr{P}}^{\mathsf{sign}})^{\mathcal{S}}$. Hence, the categorical projected Bellman equation $\bm{\eta}={\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}\bm{\eta}$ admits a unique solution $\bm{\eta}^{\pi,K}$, which satisfies $\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})\leq(1-\gamma)^{-1/2}\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},{\bm{\Pi}}_{K}\bm{\eta}^{\pi})$ (Proposition 3 Rowland et al., 2018). Applying LSA to solving the equation yields categorical TD learning, and the iteration rule is given by
$\eta_{t}(s_{t})=\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-{\bm{\Pi}}_{K}\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})].$
@@ 3 Linear-Categorical TD Learning
3.1 Linear-Categorical Parametrization of Return Distributions
Now, we combine linear function approximation with categorical parametrization, and introduce the space of linear-categorical parametrized signed measures with total mass $1$: [^3] Compared with the parametrization in (Section 9.6 Bellemare et al., 2023), our model is identifiable because we remove redundant parameters in their model, which can be linearly represented by other parameters.
${\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}:=\left\{\bm{\eta}_{{\bm{w}}}=\left({\eta_{{\bm{w}}}(s)}\right)_{s\in{\mathcal{S}}}\colon\eta_{{\bm{w}}}(s)=\sum_{k=0}^{K}p_{k}(s;{\bm{w}})\delta_{x_{k}},{\bm{w}}\in{\mathbb{R}}^{dK}\right\},$
which is an affine subspace of $({\mathscr{P}}^{\mathsf{sign}}_{K})^{{\mathcal{S}}}$, where ${\bm{w}}=({\bm{w}}(0)^{\intercal},\cdots,{\bm{w}}(K-1)^{\intercal})^{\intercal}$ and
$p_{k}(s;{\bm{w}})=\begin{cases}{\bm{\phi}}(s)^{\intercal}{\bm{w}}(k)+\frac{1}{K+1},&\text{ for}\ k\in\left\{{0,1,\cdots K-1}\right\};\\ -{\bm{\phi}}(s)^{\intercal}\sum_{j=0}^{K-1}{\bm{w}}(j)+\frac{1}{K+1},&\text{ for}\ k=K.\end{cases}$ (6)
Again, this representation is an isometry (see Proposition C.1). In many cases, it is more convenient to work with the matrix version of the parameter ${\bm{W}}:=\left({{\bm{w}}(0),\cdots,{\bm{w}}(K-1)}\right)\in{\mathbb{R}}^{d\times K}$ instead of ${\bm{w}}=\mathsf{vec}\left({{\bm{W}}}\right)$. We abbreviate ${\bm{p}}_{\bm{\eta}_{\bm{w}}}$ as ${\bm{p}}_{{\bm{w}}}$, then by Lemma A.2, for any $s\in{\mathcal{S}}$, it holds that
${\bm{p}}_{{\bm{w}}}(s)={\bm{W}}^{\intercal}{\bm{\phi}}(s)+(K+1)^{-1}\bm{1}_{K}.$ (7)
We define the linear-categorical projection operator ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\colon({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}\to{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}$ as follow:
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}:=\operatorname*{argmin}\nolimits_{\bm{\eta}_{{\bm{w}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}}\ell_{2,\mu_{\pi}}\left({\bm{\eta},\bm{\eta}_{{\bm{w}}}}\right),\quad\forall\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}.$
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$ is in fact an orthogonal projection (see Proposition C.2), and thus is non-expansive. The following proposition characterizes ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$, whose proof can be found in Appendix C.2.
@@ 3.1 Linear-Categorical Parametrization of Return Distributions
For any $\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$, ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}$ is uniquely give by $\bm{\eta}_{\tilde{\bm{w}}}$, where
$\tilde{{\bm{w}}}=\mathsf{vec}(\tilde{\bm{W}}),\quad\tilde{\bm{W}}={\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}[{\bm{\phi}}(s)\left({{\bm{p}}_{\bm{\eta}}(s)-(K+1)^{-1}\bm{1}_{K}}\right)^{\intercal}].$
3.2 Linear-Categorical Projected Bellman Equation
Since ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in the Polish space $({\mathscr{P}}_{{\bm{\phi}},K}^{\mathsf{sign}},\ell_{2,\mu_{\pi}})$, we have the following theorem, whose proof can be found in Appendix C.3.
@@ 3.2 Linear-Categorical Projected Bellman Equation
The linear-categorical projected Bellman equation $\bm{\eta}_{{\bm{w}}}={\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{w}}}$ admits a unique solution $\bm{\eta}_{{\bm{w}}^{\star}}$, where the matrix version of the parameter ${\bm{W}}^{\star}\in{\mathbb{R}}^{d\times K}$ is the unique solution to the linear system for ${\bm{W}}\in{\mathbb{R}}^{d\times K}$
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{W}}-{\mathbb{E}}_{s,s^{\prime},r}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\intercal}{\bm{W}}\tilde{{\bm{G}}}^{\intercal}(r)}\right]=\frac{1}{K+1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r)-\bm{1}_{K}}\right)^{\intercal}}\right],$ (8)
where for any $r\in[0,1]$ and $j,k\in\left\{{0,1\cdots,K}\right\}$,

### v2

@@ 2.1 Policy Evaluation
One can check that the linear projected Bellman operator ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ is a $\gamma$-contraction in the Polish space $({\mathscr{V}}_{{\bm{\phi}}},\left\|{\cdot}\right\|_{\mu_{\pi}})$. Hence, ${\bm{\Pi}}_{{\bm{\phi}}}^{\pi}\bm{T}^{\pi}$ admits a unique fixed point ${\bm{V}}_{{\bm{\psi}}^{\star}}$, which satisfies $\|{\bm{V}}^{\pi}{-}{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}\leq(1{-}\gamma^{2})^{-1/2}\|{\bm{V}}^{\pi}{-}{\bm{\Pi}}_{{\bm{\phi}}}^{\pi}{\bm{V}}^{\pi}\|_{\mu_{\pi}}$ [4, Theorem 9.8]. In Appendix C.1, we show that ${\bm{\psi}}^{\star}\in{\mathbb{R}}^{d}$ is the unique solution to the linear system for ${\bm{\psi}}\in{\mathbb{R}}^{d}$:
$\left({{\bm{\Sigma}}_{{\bm{\phi}}}-\gamma{\mathbb{E}}_{s,s^{\prime}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}}\right]}\right){\bm{\psi}}={\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)r}\right],\quad{\bm{\Sigma}}_{{\bm{\phi}}}:={\mathbb{E}}_{s\sim\mu_{\pi}}\left[{{\bm{\phi}}(s){\bm{\phi}}(s)^{\top}}\right].$ (4)
@@
In the subscript of the expectation, we abbreviate $s\sim\mu_{\pi}(\cdot),a\sim\pi(\cdot|s),(r,s^{\prime})\sim{\mathcal{P}}(\cdot,\cdot\mid s,a)$ as $s,a,r,s^{\prime}$. For brevity, we will use such abbreviations in this paper when there is no ambiguity. We can use LSA to solve the linear projected Bellman equation (Eqn. 4). As a result, at the $t$-th time-step, the updating scheme of Linear-TD is
@@
$\text{{{{Linear-TD}}}: }\qquad{\bm{\psi}}_{t}\leftarrow{\bm{\psi}}_{t-1}-\alpha{\bm{\phi}}(s_{t})\left[{\left({{\bm{\phi}}(s_{t})-\gamma{\bm{\phi}}(s_{t+1})}\right)^{\top}{\bm{\psi}}_{t-1}-r_{t}}\right].$ (5)
2.2 Distributional Policy Evaluation
In certain applications, we are not only interested in finding the expectation of random return $G^{\pi}(s)$ but also want to find the whole distribution of $G^{\pi}(s)$. This task is called distributional policy evaluation. We use $\eta^{\pi}(s)\in{\mathscr{P}}$ to denote the distribution of $G^{\pi}(s)$ and let ${\bm{\eta}}^{\pi}:=(\eta^{\pi}(s))_{s\in{\mathcal{S}}}\in{\mathscr{P}}^{\mathcal{S}}$. Then ${\bm{\eta}}^{\pi}$ satisfies the distributional Bellman equation:
$\eta^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),(r,s^{\prime})\sim{\mathcal{P}}(\cdot,\cdot\mid s,a)}[\left({b_{r,\gamma}}\right)_{\#}\eta^{\pi}(s^{\prime})],\quad\forall s\in{\mathcal{S}},$ (6)
where the RHS is the distribution of $r_{0}+\gamma G^{\pi}(s_{1})$ conditioned on $s_{0}=s$. Here $b_{r,\gamma}(x):=r+\gamma x$ for any $x\in{\mathbb{R}}$, and $f_{\#}\nu\in{\mathscr{P}}$ is defined as $f_{\#}\nu(A):=\nu(\{x\colon f(x)\in A\})$ for any function $f\colon{\mathbb{R}}\to{\mathbb{R}}$, probability measure $\nu\in{\mathscr{P}}$ and Borel set $A\subset{\mathbb{R}}$. The distributional Bellman equation can also be written as ${\bm{\eta}}^{\pi}={\bm{\bm{\mathcal{T}}}}^{\pi}{\bm{\eta}}^{\pi}$. The operator ${{\bm{\mathcal{T}}}}^{\pi}\colon{\mathscr{P}}^{\mathcal{S}}\to{\mathscr{P}}^{\mathcal{S}}$ is called the distributional Bellman operator. In this task, our goal is to find ${\bm{\eta}}^{\pi}$ for some given policy $\pi$.
Tabular Distributional TD Learning.
In analogy to tabular TD learning (Eqn. (2)), in the tabular setting, we can solve the distributional Bellman equation by LSA and derive the distributional TD learning rule given the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$:
$\eta_{t}(s_{t})\leftarrow\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})],\quad\eta_{t}(s)\leftarrow\eta_{t-1}(s),\ \forall s\neq s_{t}.$
@@ 2.2 Distributional Policy Evaluation
We comment the algorithm above is not computationally feasible as we need to manipulate infinite-dimensional objects (return distributions) at each iteration.
Categorical Parametrization and Tabular Categorical TD Learning.
@@
In order to deal with return distributions in a computationally tractable manner, we consider the categorical parametrization as in Bellemare et al. [2], Rowland et al. [48, 51], Peng et al. [43]. To be compatible with linear function approximation introduced in the next section, which cannot guarantee non-negative outputs, we will work with ${\mathscr{P}}^{\mathsf{sign}}$, the signed measure space with total mass $1$ as in Bellemare et al. [3], Lyle et al. [32], Bellemare et al. [4] instead of standard probability space ${\mathscr{P}}\subset{\mathscr{P}}^{\mathsf{sign}}$:
@@
${\mathscr{P}}^{\mathsf{sign}}:=\left\{\nu\colon\nu({\mathbb{R}})=1,\operatorname{supp}(\nu)\subseteq\left[0,(1-\gamma)^{-1}\right]\right\}.$
@@
For any $\nu\in{\mathscr{P}}^{\mathsf{sign}}$, we define its cumulative distribution function (CDF) as $F_{\nu}(x):=\nu([0,x])$. We can naturally define the $L^{2}$ and $L^{1}$ distances between CDFs as the Cramér distance $\ell_{2}$ and $1$-Wasserstein distance $W_{1}$ in ${\mathscr{P}}^{\mathsf{sign}}$, respectively. The distributional Bellman operator (see Eqn. (6)) can also be extended to the product space $({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$ without modifying its definition.
@@
${\mathscr{P}}^{\mathsf{sign}}_{K}:=\left\{\nu_{\bm{p}}=\sum_{k=0}^{K}p_{k}\delta_{x_{k}}\colon{\bm{p}}=\left({p_{0},\ldots,p_{K{-}1}}\right)^{\top}\in{\mathbb{R}}^{K},p_{K}=1-\sum_{k=0}^{K-1}p_{k}\right\},$ (7)
which is an affine subspace of ${\mathscr{P}}^{\mathsf{sign}}$. Here $\left\{{x_{k}{=}k\iota_{K}}\right\}_{k{=}0}^{K}$ are $K{+}1$ equally-spaced points of the support, $\iota_{K}{=}\left[{K(1{-}\gamma)}\right]^{{-}1}$ is the gap between adjacent points, and $p_{k}$ is the ‘probability’ (may be negative) that $\nu$ assigns to $x_{k}$. We define the categorical projection operator ${\bm{\Pi}}_{K}{\colon}{\mathscr{P}}^{\mathsf{sign}}{\to}{\mathscr{P}}^{\mathsf{sign}}_{K}$ as
@@
Following Bellemare et al. [4, Proposition 5.14], one can show that $\bm{\Pi}_{K}\nu\in{\mathscr{P}}^{\mathsf{sign}}_{K}$ is uniquely represented with a vector ${\bm{p}}_{\nu}=\left({p_{k}(\nu)}\right)_{k=0}^{K-1}\in{\mathbb{R}}^{K}$, where
$p_{k}(\nu)=\int_{\left[{0,(1-\gamma)^{-1}}\right]}(1-\left|{(x-x_{k})/{\iota_{K}}}\right|)_{+}\nu(dx).$ (8)
We lift ${\bm{\Pi}}_{K}$ to the product space by defining $\left[{\bm{\Pi}_{K}{\bm{\eta}}}\right](s):=\bm{\Pi}_{K}\eta(s)$. One can check that the categorical Bellman operator ${\bm{\Pi}}_{K}{\bm{\mathcal{T}}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in the Polish space $(({\mathscr{P}}_{K}^{\mathsf{sign}})^{\mathcal{S}},\ell_{2,\mu_{\pi}})$, where $\ell_{2,\mu_{\pi}}(\bm{\eta}_{1},\bm{\eta}_{2}):=({\mathbb{E}}_{s\sim\mu_{\pi}}[\ell_{2}^{2}(\eta_{1}(s),\eta_{2}(s))])^{1/2}$ is the $\mu_{\pi}$-weighted Cramér distance between $\bm{\eta}_{1},\bm{\eta}_{2}\in({\mathscr{P}}^{\mathsf{sign}})^{\mathcal{S}}$. Similarly, $W_{1,\mu_{\pi}}(\bm{\eta}_{1},\bm{\eta}_{2}):=({\mathbb{E}}_{s\sim\mu_{\pi}}[W_{1}^{2}(\eta_{1}(s),\eta_{2}(s))])^{1/2}$. Hence, the categorical projected Bellman equation $\bm{\eta}={\bm{\Pi}}_{K}\bm{\mathcal{T}}^{\pi}\bm{\eta}$ admits a unique solution $\bm{\eta}^{\pi,K}$, which satisfies $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})\leq(1{-}\gamma)^{-1}\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},{\bm{\Pi}}_{K}\bm{\eta}^{\pi})$ [48, Proposition 3]. Applying LSA to solving the equation yields tabular categorical TD learning, and the iteration rule is given by
$\eta_{t}(s_{t})\leftarrow\eta_{t-1}(s_{t})-\alpha[\eta_{t-1}(s_{t})-{\bm{\Pi}}_{K}\left({b_{r_{t},\gamma}}\right)_{\#}\eta_{t-1}(s_{t+1})],\quad\eta_{t}(s)\leftarrow\eta_{t-1}(s),\ \forall s\neq s_{t}.$ (9)
@@ 3 Linear-Categorical TD Learning
In this section, we propose our Linear-CTD algorithm (Eqn. (13)) by combining the linear function approximation (Eqn. (3)) with the categorical parametrization (Eqn. (7)). We first introduce the space of linear-categorical parametrized signed measures with total mass $1$:
${\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}:=\left\{\bm{\eta}_{{\bm{\theta}}}=\left({\eta_{{\bm{\theta}}}(s)}\right)_{s\in{\mathcal{S}}}\colon\eta_{{\bm{\theta}}}(s)=\sum_{k=0}^{K}p_{k}(s;{\bm{\theta}})\delta_{x_{k}},{\bm{\theta}}=({\bm{\theta}}(0)^{\top},\ldots,{\bm{\theta}}(K{-}1)^{\top})^{\top}\in{\mathbb{R}}^{dK}\right\},$
which is an affine subspace of $({\mathscr{P}}^{\mathsf{sign}}_{K})^{{\mathcal{S}}}$. Here $p_{k}(s;{\bm{\theta}}){=}F_{k}(s;{\bm{\theta}}){-}F_{k{-}1}(s;{\bm{\theta}})$, and
$F_{k}(s;{\bm{\theta}})={\bm{\phi}}(s)^{\top}{\bm{\theta}}(k)+({k+1})/({K+1})\quad\text{ for}\ k\in\left\{{0,1,\ldots,K-1}\right\}$ (10)
[…]
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}:=\operatorname*{argmin}\nolimits_{\bm{\eta}_{{\bm{\theta}}}\in{\mathscr{P}}^{\mathsf{sign}}_{{\bm{\phi}},K}}\ell_{2,\mu_{\pi}}\left({\bm{\eta},\bm{\eta}_{{\bm{\theta}}}}\right),\quad\forall\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}.$
${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$ is in fact an orthogonal projection (see Proposition D.2), and thus is non-expansive. The following proposition characterizes ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}$, whose proof can be found in Appendix D.2.
@@
For any $\bm{\eta}\in({\mathscr{P}}^{\mathsf{sign}})^{{\mathcal{S}}}$, ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}$ is uniquely given by $\bm{\eta}_{\tilde{\bm{\theta}}}$, where $\tilde{{\bm{\theta}}}=\mathsf{vec}(\tilde{\bm{\Theta}})$,
$\tilde{\bm{\Theta}}{=}{\bm{\Sigma}}_{{\bm{\phi}}}^{-1}{\mathbb{E}}_{s\sim\mu_{\pi}}[{\bm{\phi}}(s)({\bm{p}}_{\bm{\eta}}(s){-}(K{+}1)^{-1}\bm{1}_{K})^{\top}{\bm{C}}^{\top}],\quad{\bm{C}}{=}\left[{\mathds{1}\left\{{i\geq j}\right\}}\right]_{i,j\in[K]}\in{\mathbb{R}}^{K{\times}K}.$ (11)
Here ${\bm{p}}_{\bm{\eta}}(s){:}{=}{\bm{p}}_{\eta(s)}{=}(p_{k}(\eta(s)))_{k{=}0}^{K{-}1}$ is the vector that identifies ${\bm{\Pi}}_{K}\eta(s)$ defined in Eqn. (8).
Since ${\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}$ is a $\sqrt{\gamma}$-contraction in $({\mathscr{P}}_{{\bm{\phi}},K}^{\mathsf{sign}},\ell_{2,\mu_{\pi}})$ ($\bm{\mathcal{T}}^{\pi}$ is $\sqrt{\gamma}$-contraction [4, Lemma 9.14]), in the following theorem, we can generalize the linear projected Bellman equation (Eqn. (4)) to the distributional setting. The proof can be found in Appendix D.3.
@@
The linear-categorical projected Bellman equation $\bm{\eta}_{{\bm{\theta}}}{=}{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\mathcal{T}}^{\pi}\bm{\eta}_{{\bm{\theta}}}$ admits a unique solution $\bm{\eta}_{{\bm{\theta}}^{\star}}$, where the matrix parameter ${\bm{\Theta}}^{\star}$ is the unique solution to the linear system for ${\bm{\Theta}}{\in}{\mathbb{R}}^{d{\times}K}$
${\bm{\Sigma}}_{{\bm{\phi}}}{\bm{\Theta}}{-}{\mathbb{E}}_{s,s^{\prime},r}\left[{{\bm{\phi}}(s){\bm{\phi}}(s^{\prime})^{\top}{\bm{\Theta}}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{{-}1})^{\top}}\right]{=}\frac{1}{K{+}1}{\mathbb{E}}_{s,r}\left[{{\bm{\phi}}(s)(\sum_{j=0}^{K}{\bm{g}}_{j}(r){-}\bm{1}_{K})^{\top}{\bm{C}}^{\top}}\right],$ (12)
where for any $r\in[0,1]$ and $j,k\in\left\{{0,1,\ldots,K}\right\}$,